OpenStudy (anonymous):

If \[\lim_{h \rightarrow 0} \frac{ \arcsin(a+h)-\arcsin(a) }{ h } = 2\] , which of the following could be the value of a? A) sqrt2/2 B) sqrt3/2 C)sqrt3 D)1/2 E) 2

OpenStudy (anonymous):

The left side is straight from the definition of the derivative: \[\lim_{h\to0}\frac{f(x+h)-f(x)}{h}=:f'(x)\] As for values of the derivative, you have \[\lim_{x\to c}\frac{f(x)-f(c)}{x-c}=:f'(c)\] so you have \(f(x)=\arcsin x\) and (\(c=a\) in this problem). Recall the derivative of the inverse sine function: \[f(x)=\arcsin x~~\implies~~f'(x)=\frac{1}{\sqrt{1-x^2}}\] and so given that \(f'(a)=2\), you have \[\frac{1}{\sqrt{1-a^2}}=2\]